# Proximal Supervised Fine-Tuning

### v1

@@ Abstract
Supervised fine-tuning (SFT) of foundation models often leads to poor generalization, where prior capabilities deteriorate after tuning on new tasks or domains. Inspired by trust-region policy optimization (TRPO) and proximal policy optimization (PPO) in reinforcement learning (RL), we propose Proximal SFT (PSFT), a fine-tuning objective that incorporates the benefits of trust-region, effectively constraining policy drift during SFT while maintaining competitive tuning. By viewing SFT as a special case of policy gradient methods with constant positive advantages, we derive PSFT that stabilizes optimization and leads to generalization, while leaving room for further optimization in subsequent post-training stages. Experiments across mathematical and human-value domains show that PSFT matches SFT in-domain, outperforms it in out-of-domain generalization, remains stable under prolonged training without causing entropy collapse, and provides a stronger foundation for the subsequent optimization.111Code is available at https://github.com/zwhong714/PSFT.
@@ 1 Introduction
Recently, post-training has become a crucial part of the overall training process. In particular, reinforcement learning (RL) algorithms, such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024), have demonstrated significant effectiveness when applied to language models (LMs) focused on reasoning tasks. As RL is scaled over time, foundation models gain the capacity to address complex problems through more profound and extended reasoning (OpenAI, 2024; Guo et al., 2025). These reasoning models offer an abundant and valuable latent thoughts (Ruan et al., 2025) across the internet. Numerous community efforts have focused on leveraging this knowledge via supervised fine-tuning (SFT) (Guha et al., 2025; Li et al., 2025), a distillation approach valued for its efficiency and simplicity compared to RL.
@@
Another concern is maintaining the ability to explore. In practice, SFT is often used as a cold start to stabilize RFT training. The gains from RFT may largely come from refining capabilities already acquired during pretraining and SFT (Gandhi et al., 2025). However, excessive reliance on SFT can diminish a model’s capacity for exploration (Xie et al., 2024), as it would cause the entropy collapse (Cui et al., 2025) and thereby constrain exploration during RL training (Yu et al., 2025).
Therefore, the current challenges lie in improving the generalization and exploration keeping of SFT models. To tackle these challenges, this paper proposes an improved fine-tuning strategy Proximal Supervised Fine-Tuning (PSFT) that avoids rote learning and achieves reliable performance. Specifically, we establish a theoretical connection between SFT and RL, and introduce a novel objective based on a clipped surrogate objective that leverages the benefits of trust regions to constrain policy updates.
Our experiments demonstrate that, compared to standard SFT, PSFT attains comparable performance on the target task while preserving the model’s general capabilities. Moreover, PSFT mitigates entropy collapse during training and yields superior target and generalization performance in subsequent RL stages. We evaluate PSFT on both the mathematics and human value alignment domains, highlighting the broad applicability of PSFT in practice.
Our contributions are as follows:
We propose PSFT, an optimization method for SFT that adopts a clipped surrogate objective, akin to PPO, to enforce trust-region-like constraints and thereby prevent excessive policy update. This approach maximally preserves the model’s general capabilities while maintaining comparable performance on the target task.
PSFT effectively prevents entropy collapse and overfitting during SFT, thereby enabling subsequent RL stages to achieve more robust and superior results on both target-specific and general tasks.
We extensively validate PSFT across various base models, tasks, and evaluation metrics, demonstrating its potential as a promising alternative to standard SFT.
@@ 2 Preliminaries and Motivations
2.1 Markov Decision Process Formulation
We formulate the problem of language modeling and optimization in the framework of a Markov Decision Process (MDP). An MDP is defined by a tuple , where: is the state space, representing partial sequences. is the action space, representing possible next tokens. is the transition probability of moving from state to given action . In the context of auto-regressive LMs, a query sequence initializes the state. A response sequence is generated by sequentially sampling actions (tokens) from the policy . The joint probability of given is: , where and the state at time is .
@@ 2.2 SFT as the special policy update
Supervised Fine-tuning.
The training objective is to minimize the cross-entropy loss between the model’s predicted token distribution and the ground truth tokens. Formally, the loss is defined as:
where pairs are sampled from an offline dataset .
Policy Gradient.
In contrast, policy gradient directly samples trajectories from the current policy interacting with the environment. Using the policy gradient theorem, the corresponding loss function can be defined as follows:
where is the estimated advantage function at step .
From this perspective, SFT can be seen as a special case of policy gradient where sampling is from a fixed offline dataset and the advantage is fixed as for the ground-truth action, which corresponds to maximizing the likelihood of expert tokens.
2.3 Proximal Policy Optimization
Trust Region Policy Optimization.
TRPO (Schulman et al., 2015) maximizes a surrogate objective by leveraging trajectories collected under the previous policy , introducing importance sampling to reweight these samples for evaluating the new policy :
where is the importance sampling ratio, and the superscript CPI refers to conservative policy iteration (Kakade & Langford, 2002). By introducing a trust region constraint on the KL divergence between and , TRPO ensures that each policy update does not deviate too far from the reference policy.
Proximal Policy Optimization.
However, directly maximizing with a hard KL constraint, as in TRPO, can be difficult to optimize in practice. To address this, PPO (Schulman et al., 2017) modifies the surrogate objective by penalizing policy changes that move too far from 1. The clipped surrogate objective is:
This clipping mechanism effectively defines a soft trust region: is restricted to remain within a small neighborhood of 1, which prevents destructive updates. Compared to TRPO, PPO achieves similar stability and efficiency.
@@ 3 Proximal Supervised Fine-Tuning
Proximal Supervised Fine-Tuning aims to improve performance on the target SFT task while preserving general capabilities, preventing entropy collapse, and allowing further optimization.
3.1 From PPO to PSFT
We revisit the TRPO and PPO objectives in RL. Both methods rely on importance sampling ratios to reweight returns, while constraining these ratios to prevent the new policy from deviating excessively from the old policy . Translating this idea to the supervised setting, where all actions are assumed to be “correct” (i.e., ), we simplify the advantage to , and define the Proximal SFT loss as:
This objective regularizes supervised learning updates by limiting the ratio between the new and old policy probabilities. We allow the old policy —rather than the fixed initial model—to evolve dynamically. Intuitively, it discourages overconfident changes in token probabilities, thereby preserving existing capabilities. When changing to the online setting, this method is equivalent to standard SFT training.
Warm-Up.
Note that in Eq. 5 is sampled from the offline dataset . In the initial steps, may yield a biased expectation since is not aligned with the distribution of . The clipping mechanism constrains policy updates within an acceptable range, which may indirectly reduce the negative impact of this mismatch. A warm-up SFT phase on can be introduced to better align the initial policy with the offline dataset, further improving the SFT/PSFT in-domain performance.
3.2 Gradient Analysis
The gradient updated is confined to the trust region to perform fine-tuning, similar to that in PPO. The gradient is as follows:
The gradient analysis shows that if the train dataset distribution deviates significantly from the model distribution, these tokens will have no gradient, thereby avoiding large policy updates and maintaining generalization. Typically, the optimal value of is set to or . A larger can lead to large gradients, as shown in Equation 6. Further analysis is provided in Section 5.2.
@@ 4 Experiments
In Section 4.1, we explore the training dynamics of our PSFT with standard SFT, and then evaluate both in-domain and out-of-domain performances of models fine-tuned with each method, focusing primarily on the mathematics domain. In Section 4.2, we attempt to verify the effectiveness of PSFT as the cold start point of RL. In Section 4.3, we further conduct experiments in the alignment domain to assess the universal applicability of our approach.
@@ 4.1 Main Experiments on Math Reasoning in the SFT stage
(1) Models and Datasets. We evaluate our method based on Qwen2.5-7B-Instruct (Yang et al., 2025) and Llama3.1-8B-Instruct (Dubey et al., 2024). Our training data focuses on the math domain, aiming to improve general LLM capabilities through math reasoning. We employ the OpenR1-Math-8192222The dataset undergoes filtering through Math-Verify to exclude instances with unverifiable answers or responses exceeding 8,192 tokens in length. long chain-of-thought (CoT) dataset (Face, 2025). (2) Baseline. We consider the SFT method and an SFT variant, denoted as , which incorporates KL divergence constraints into the loss function. The KL regularization coefficient is set to 0.5, while the parameter in PSFT is fixed at . More details can be found in Appendix A.1.1.
4.1.1 Training Dynamics
@@
We begin by examining the training dynamics of each method. Specifically, we train each for around 10 epochs and plot the corresponding changes in entropy and performance. In this section, we use AIME-24 avg@32 to measure in-domain performance and GPQA (Rein et al., 2024) avg@8 to assess out-of-domain performance.
Observation 1: PSFT avoids entropy collapse.
@@
The entropy evolution is shown in Figure 1. Compared to SFT and SFT-KL, PSFT produces a smoother entropy curve. For SFT and SFT-KL, entropy exhibits a marked decline after each epoch, indicating potential overfitting. This suggests that PSFT is capable of sustaining long-term, token-level fine-grained training without triggering entropy collapse. Notably, PSFT with a warm-up phase demonstrates the same stability, and the number of warm-up rounds plays a critical role in shaping the overall entropy level. Further evidence is provided in Figure 6(a).
Analysis 1: PSFT matches or surpasses in-domain performance of standard SFT at a similar entropy level.
@@
The in-domain evaluation performance evolution is shown in Figure 2. PSFT achieves comparable performance on targeted math tasks and can even outperform SFT under the same entropy level. For instance, at step 1300, PSFT attains an AIME24 score of nearly 20 with an entropy around 0.3. This entropy level corresponds to that of SFT between epochs 300 and 520, yet PSFT achieves a higher AIME score than SFT within this range. While the KL-based approach intuitively addresses distribution shift, considering both entropy trends and in-domain performance, it remains less effective than PSFT.
Analysis 2: PSFT with warm-up operation surpasses standard SFT.
Figure 2 illustrates the impact of introducing a warm-up phase into PSFT. We observe that the in-domain performance increases steadily. Across both models, PSFT with warm-up consistently outperforms vanilla PSFT as well as the standard SFT baselines. Moreover, increasing the number of warm-up steps further enhances in-domain performance. This offers an effective way to fully exploit high-quality datasets such as s1k (Muennighoff et al., 2025) and LIMO (Ye et al., 2025) while avoiding entropy collapse.
Analysis 3: PSFT can largely maintain the generalization of the model.
The evolution of out-of-domain evaluation performance is shown in Figure 3. Increasing the number of training steps leads to a decline in generalization for SFT models. For example, after 700 steps, the SFT method on Qwen2.5-7B-Instruct exhibits a clear decline in performance. While SFT-KL alleviates this degradation to a certain degree, PSFT demonstrates substantially better generalization.
@@
Overall, benefiting from its robust entropy control during training, PSFT avoids large fluctuations in policy updates. As observed in both in-domain and out-of-domain evaluations, PSFT generally maintains an upward performance trend. In contrast, SFT and SFT-KL exhibit pronounced fluctuations in out-of-domain performance.
@@ 4.1.2 Detailed Evaluations
(1) Model setup: For a fair comparison, we select checkpoints based on in-domain performance. For instance, on Qwen2.5-7B-Instruct, we use the SFT model trained for 700 steps, the SFT-KL model trained for 900 steps, and the PSFT model trained for 1300 steps. (2) Evaluation benchmark: (i) In-domain tasks: AIME24, AIME25, AMC, MATH-500 (Hendrycks et al., 2021), OlympidBench (He et al., 2024), and Minerva (Lewkowycz et al., 2022). (ii) Out-of-domain tasks: GPQA (Rein et al., 2024), ARC-C (Clark et al., 2018), TruthfulQA (Lin et al., 2021), MMLU-Pro (Wang et al., 2024), SuperGPQA (Du et al., 2025), HeadQA (Vilares & Gómez-Rodríguez, 2019) and IFEval (Zhou et al., 2023). (3) Inference. The inference length is set to 10,240 tokens (4,096 for IFEval), with a top- of 0.95 and a temperature of 0.7.
| Method | AIME24 | AIME25 | AMC | MATH-500 | OlympidBench | Minerva | Avg. |
| Qwen2.5-7B-Instruct | |||||||
| Original | 11.25 | 8.75 | 52.58 | 75.05 | 39.72 | 40.53 | 37.98 |
| SFT | 22.08 | 23.02 | 62.73 | 84.10 | 52.35 | 43.66 | 47.99 |
| SFT | 19.27 | 21.56 | 63.20 | 83.55 | 52.70 | 42.19 | 47.08 |
[…]
| Llama3.1-8B-Instruct | |||||||
| Original | 3.96 | 0.73 | 24.45 | 48.55 | 17.09 | 25.87 | 20.11 |
| SFT | 10.63 | 16.77 | 47.19 | 72.60 | 41.43 | 32.17 | 36.80 |
| SFT | 9.58 | 16.15 | 45.55 | 69.83 | 39.19 | 26.75 | 34.51 |
[…]
Analysis 1. PSFT with warm-up consistently outperforms standard SFT on in-domain tasks.
@@
Table 1 presents the performance across various in-domain benchmarks, where all fine-tuned methods achieve notable improvements over the original model. PSFT with a warm-up phase not only stabilizes early training dynamics as illustrated in Figure 2, but also enhances robustness across different evaluation benchmarks.
| Method | GPQA | ARC-C | TruthfulQA | MMLU-Pro | SuperGPQA | HeadQA | IFEval | Avg. |
| Qwen2.5-7B-Instruct | ||||||||
| Original | 31.38 | 91.54 | 66.10 | 54.99 | 27.59 | 73.41 | 73.94 | 59.85 |
| SFT | 32.89 | 92.22 | 63.14 | 58.98 | 29.02 | 74.65 | 54.42 | 57.90 |
| SFT | 32.95 | 91.86 | 61.31 | 58.35 | 27.69 | 74.07 | 55.44 | 57.38 |
[…]
| LLama3.1-8B-Instruct | ||||||||
| Original | 24.62 | 80.96 | 55.14 | 43.70 | 18.16 | 68.20 | 78.10 | 52.70 |
| SFT | 19.38 | 87.73 | 67.08 | 50.29 | 21.95 | 73.52 | 33.45 | 50.49 |
| SFT | 18.18 | 87.02 | 67.03 | 47.02 | 19.81 | 71.95 | 34.19 | 49.31 |
[…]
Analysis 2: PSFT demonstrates strong generalization.
@@
Table 2 shows different performance on out-of-domain benchmarks. Overall, our results are consistent with those reported in Zhou et al. (2025): injecting long CoT data into the model improves its general reasoning ability. However, on IFEval, which evaluates instruction-following capability, SFT, SFT-KL, and PSFT with warm-up significantly degrade the original performance. In contrast, PSFT not only preserves instruction compliance but also substantially enhances the model’s reasoning ability across other domains. PSFT with a warm-up phase also consistently surpasses standard SFT on out-of-domain tasks.
Analysis 3: PSFT demonstrates robustness across models.
@@
For example, when evaluating GPQA and TruthfulQA tasks, SFT and SFT-KL exhibit varying behaviors across different models—one sometimes compromises performance while the other may boost it. In contrast, PSFT consistently improves performance, demonstrating a strong advantage over Llama models.

### v2

@@ Abstract
Supervised fine-tuning (SFT) of foundation models often leads to poor generalization, where prior capabilities deteriorate after tuning on specific tasks. Inspired by trust-region policy optimization (TRPO) and proximal policy optimization (PPO) in reinforcement learning (RL), we propose Proximal SFT (PSFT), a fine-tuning objective that incorporates the benefits of trust-region, effectively constraining policy drift during SFT while maintaining competitive tuning. By viewing SFT as a special case of policy gradient methods with constant positive advantages, we derive PSFT that stabilizes optimization and leads to generalization, while leaving room for further optimization in subsequent post-training stages. Experiments across mathematical, human-value, and multimodal domains show that PSFT matches standard SFT in-domain, outperforms it in out-of-domain generalization, remains stable under prolonged training without causing entropy collapse, and provides a stronger foundation for the subsequent optimization.
@@ 1 Introduction
Recently, post-training has become a crucial part of the overall training process. In particular, reinforcement learning (RL) algorithms, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have demonstrated significant effectiveness when applied to language models (LMs) focused on reasoning tasks. As RL is scaled over time, foundation models gain the capacity to address complex problems through more profound and extended reasoning (OpenAI, 2024; Guo et al., 2025; Zhu et al., 2025), producing high-quality reasoning trajectories. Numerous community efforts have focused on leveraging this knowledge via supervised fine-tuning (SFT) (Guha et al., 2025; Li et al., 2025), a distillation approach valued for its efficiency and simplicity compared to RL.
@@
Another concern is maintaining the ability to explore. In practice, SFT is often used as a cold start to stabilize RFT training. The gains from RFT may largely come from refining capabilities already acquired during pre-training and SFT (Gandhi et al., 2025). However, excessive reliance on SFT can diminish a model’s capacity for exploration (Xie et al., 2024), as it would cause the entropy collapse (Cui et al., 2025) and thereby constrain exploration during RL training (Yu et al., 2025).
Therefore, the current challenges lie in improving the generalization and exploration abilities of SFT models. To tackle these challenges, this paper proposes an improved fine-tuning strategy Proximal Supervised Fine-Tuning (PSFT) that avoids rote learning and achieves reliable performance in post-training. Specifically, we establish a theoretical connection between SFT and RL, and then introduce a novel objective based on a clipped surrogate objective from PPO to SFT, which leverages the benefits of trust regions to constrain policy updates.
Our experiments demonstrate that, compared to standard SFT, PSFT attains comparable performance on the target task while preserving the model’s general capabilities. Moreover, PSFT mitigates entropy collapse during training and yields superior target and generalization performance in subsequent RL stages. We evaluate PSFT on mathematics, human value alignment, and multimodal tasks, highlighting its broad applicability in practice. Our contributions are as follows:
We propose PSFT, a novel SFT optimization method that utilizes a clipped surrogate objective to enforce trust-region-like constraints, preventing excessive policy updates. Compared to standard SFT, it preserves the model’s general capabilities while achieving comparable performance on target tasks.
PSFT effectively prevents entropy collapse and overfitting during SFT, thereby enabling subsequent RL stages to achieve more robust and superior results on both target and general tasks.
We extensively validate PSFT across various base models, tasks, settings, and evaluation metrics, demonstrating its potential as a promising alternative to standard SFT.
@@ 2 Preliminaries and Motivations
We formulate autoregressive language modeling as a Markov Decision Process (MDP) . Here, denotes the state space ( the prefix of the current decoding step), the action space (possible next tokens), and the transition probability , and the reward function. Given a query , the initial state is . A response is generated by sequentially selecting actions . The policy defines the probability of given as
@@ 2.1 SFT as the special policy update
Supervised Fine-tuning. The training objective of SFT is to minimize the cross-entropy loss between the model’s predicted token distribution and the ground truth tokens, which is defined as:
where pairs are sampled from an offline dataset , with representing the ground-truth token.
Policy Gradient. In contrast, policy gradient directly samples trajectories from the current policy as it interacts with the environment. Using the policy gradient theorem (Sutton et al., 2000), the objective is to maximize the expected return:
where is the estimated advantage function at step , which measures how much better an action is compared to the average action at state under the current policy.
From this perspective, SFT can be seen as a special case of policy gradient where the sampling is from a fixed offline dataset , and the advantage is fixed as for the ground-truth actions, which corresponds to maximizing the likelihood of ground-truth tokens.
2.2 Trust-Region Constraints in RL
Trust Region Policy Optimization. TRPO (Schulman et al., 2015) maximizes a surrogate objective by leveraging trajectories collected under the old policy , introducing an importance sampling ratio to reweight these samples for evaluating the new policy :
where the superscript CPI refers to conservative policy iteration (Kakade and Langford, 2002). TRPO ensures that each policy update stays within a trust region by enforcing a hard KL constraint, preventing the new policy from deviating too far from the previous one and thus maintaining stable and reliable learning.
Proximal Policy Optimization. However, directly maximizing with a hard KL constraint, as in TRPO, can be difficult to optimize in practice because of complex second-order optimization and sensitivity to the constraint threshold (Schulman et al., 2017). To address this, PPO (Schulman et al., 2017) modifies the surrogate objective by clipping the importance sampling ratio, formalized as:
This clipping mechanism effectively defines a soft trust region , which is restricted to remain within a small neighborhood of 1 (controlled by ) to prevent destructive updates. Compared to TRPO, PPO achieves greater stability and efficiency.
@@ 3 Proximal Supervised Fine-Tuning
PSFT aims to improve performance on the target task while preserving general capabilities beyond the target domain—such as scientific reasoning and instruction following—and preventing entropy collapse (the over-concentration of the output distribution that reduces diversity and harms generalization (Ziegler et al., 2019)), thereby enabling further optimization in the RL stage.
PSFT. We revisit the TRPO and PPO objectives in RL. Both methods rely on importance sampling ratios to reweight returns, while constraining these ratios to prevent the new policy from deviating excessively from the old policy . Translating this idea to the supervised setting with the offline dataset , where all actions are assumed to be “correct” (i.e., ), we simplify the advantage to (as in SFT), and define our PSFT loss as follows:
This objective regularizes supervised learning updates by limiting the ratio between the new and old policy probabilities. Intuitively, it discourages overconfident changes in token probabilities, thereby optimizing target tasks while preserving existing general capabilities.
Comparison with PPO. Although PSFT adopts the soft trust region mechanism used in PPO, it is not a policy-gradient RL method. PPO optimizes expected return using on-policy trajectories and advantage estimates; by contrast, PSFT is a supervised offline objective (here we simplify to ), where the clipped ratio functions mainly as a regularizer that limits overly large changes in token probabilities and reweights gradients.
Dynamic update of . Allowing the old policy to evolve dynamically—rather than keeping the initial model fixed—is essential for gradual and stable learning from the dataset. In contrast, fixing the initial model confines optimization to the trust region centered on the reference model, which limits the knowledge that can be effectively leveraged from the offline dataset (see Section 5.3).
Warm-Up. in Eq. 6 is sampled from the offline dataset . In the initial steps, may yield a biased expectation since may not be aligned with the distribution of . We let evolve during SFT, gradually learning from , which could largely alleviate the potential bias issue. Besides, a warm-up SFT phase on can be optionally introduced to better align the initial old policy with the offline dataset, further improving PSFT’s target-domain performance (see Section 4.1).
Gradient Analysis. Similar to PPO, the gradient is confined to the trust region for fine-tuning:
It shows that if the offline dataset distribution deviates significantly from the model distribution (which is more likely to disturb the general ability), these tokens will have no gradient. This mechanism is simple yet effective: it prevents large policy updates, mitigates entropy collapse, and preserves generalization, thereby enabling sustainable progress in the subsequent RL stage. Typically, the optimal value of is set to or (a larger can lead to large gradients). The clipped token cloud is given in Section 5.1, and further analysis is provided in Section 5.2.
@@ 4 Experiments
In Section 4.1, we explore the training dynamics and then evaluate both in-domain and out-of-domain performances. In Section 4.2, we verify the effectiveness of PSFT as the foundation of RL. In Section 4.3, we further conduct experiments in the alignment domain to assess the universal applicability of our approach. In Section 4.4, we further extend PSFT to vision language models.
@@ 4.1 Main Experiments on Math Reasoning in the SFT Stage
Setup. (1) Models and Datasets. We evaluate our method based on Qwen2.5-7B-Instruct (Yang et al., 2025) and Llama3.1-8B-Instruct (Dubey et al., 2024). Our training data focuses on the math domain, aiming to improve general LLM capabilities through math reasoning. We employ the OpenR1-Math-8192 long chain-of-thought (CoT) dataset (Face, 2025). (2) Baseline. We consider the standard SFT method and an SFT variant, denoted as , which incorporates KL divergence constraints into the loss function. The KL regularization coefficient is set to 0.5, while the parameter in PSFT is fixed at . More details can be found in Appendix C.7.1.
4.1.1 Training Dynamics on SFT
@@
We begin by examining the training dynamics of each method. Specifically, we train each for around 10 epochs and plot the corresponding changes in entropy and performance. We report in-domain performance with AIME-24 avg@32 and out-of-domain performance with GPQA avg@8.
@@
Finding 1: PSFT avoids entropy collapse. The entropy evolution is shown in Figure 1. Compared to SFT and SFT-KL, our PSFT produces a smoother entropy curve. For SFT and SFT-KL, entropy exhibits a marked decline after each epoch, indicating potential overfitting. This suggests that PSFT is capable of sustaining long-term fine-grained training without triggering entropy collapse. Notably, PSFT with a warm-up phase demonstrates the same stability, and the extent of warm-up plays a critical role in shaping the overall entropy level.
@@
Finding 2: PSFT matches or surpasses in-domain performance of standard SFT at a similar entropy level. Figure 3(a) illustrates the evolution of in-domain evaluation performance. PSFT achieves performance on par with, and sometimes exceeding, SFT under comparable entropy levels. For example, at step 1300, PSFT reaches an AIME24 score of nearly 20 with an entropy of around 0.3. This entropy corresponds to that of SFT between epochs 300 and 520, yet PSFT delivers a higher AIME score within this range. SFT with KL remains less effective than PSFT when considering both entropy dynamics and in-domain performance.
Finding 3: PSFT with warm-up surpasses standard SFT. Figure 3(a) shows the effect of adding a warm-up phase to PSFT. The in-domain performance improves steadily, with PSFT + warm-up consistently outperforming both vanilla PSFT and standard SFT baselines. Furthermore, longer warm-up periods yield even greater gains in in-domain performance.
@@
Finding 4: PSFT avoids large fluctuations in policy updates. As observed in both in-domain and out-of-domain evaluations (GPQA, Figure 3(b)), PSFT generally maintains an upward performance trend. In contrast, SFT and SFT-KL exhibit pronounced fluctuations in out-of-domain performance. Therefore, PSFT preserves the better generalization to out-of-domain tasks.
@@ 4.1.2 Detailed Evaluations
Setup. (1) Model setup: We select checkpoints by in-domain performance: for Qwen2.5-7B-Instruct, the SFT model at 700 steps, SFT-KL at 900 steps, and PSFT at 1300 steps. (2) Evaluation benchmark: See Appendix C.7.1. (3) Inference. The inference length is set to 10,240 tokens (4,096 for IFEval), with a top- of 0.95 and a temperature of 0.7.
| Qwen2.5-7B-Instruct | Llama3.1-8B-Instruct | ||||||||||
| Benchmark | Base | SFT | SFT-KL | PSFT | PSFT | Base | SFT | SFT-KL | PSFT | PSFT | |
| In-Domain | AIME-24 | 11.25 | 22.08 | 19.27 | 22.92 | 19.38 | 3.96 | 10.63 | 9.58 | 12.08 | 10.31 |
| AIME-25 | 8.75 | 23.02 | 21.56 | 23.02 | 21.98 | 0.73 | 16.77 | 16.15 | 18.75 | 14.48 | |
| AMC | 52.58 | 62.73 | 63.20 | 62.42 | 62.34 | 24.45 | 47.19 | 45.55 | 49.45 | 46.80 | |
| MATH-500 | 75.05 | 84.10 | 83.55 | 84.68 | 83.35 | 48.55 | 72.60 | 69.83 | 74.15 | 71.98 | |
| OlympidBench | 39.72 | 52.35 | 52.70 | 52.30 | 51.50 | 17.09 | 41.43 | 39.19 | 42.07 | 39.61 | |
| Minerva | 40.53 | 43.66 | 42.19 | 43.38 | 43.33 | 25.87 | 32.17 | 26.75 | 33.64 | 32.40 | |
| Avg. | 37.98 | 47.99 | 47.08 | 48.17 | 46.98 | 20.11 | 36.80 | 34.51 | 38.36 | 35.93 | |
| Out-of-Domain | GPQA | 31.38 | 32.89 | 32.95 | 33.27 | 33.21 | 24.62 | 19.38 | 18.18 | 23.99 | 26.89 |
| ARC-C | 91.54 | 92.22 | 91.86 | 92.09 | 92.29 | 80.96 | 87.73 | 87.02 | 89.72 | 89.11 | |
| TruthfulQA | 66.10 | 63.14 | 61.31 | 66.37 | 67.16 | 55.14 | 67.08 | 67.03 | 68.55 | 68.69 | |
| MMLU-Pro | 54.99 | 58.98 | 58.35 | 59.28 | 59.18 | 43.70 | 50.29 | 47.02 | 56.03 | 56.58 | |
| SuperGPQA | 27.59 | 29.02 | 27.69 | 28.37 | 28.10 | 18.16 | 21.95 | 19.81 | 25.26 | 25.85 | |
| HeadQA | 73.41 | 74.65 | 74.07 | 75.24 | 75.82 | 68.20 | 73.52 | 71.95 | 77.71 | 77.90 | |
| IFEval | 73.94 | 54.42 | 55.44 | 55.07 | 73.03 | 78.10 | 33.45 | 34.19 | 33.08 | 69.75 | |
| Avg. | 59.85 | 57.90 | 57.38 | 58.53 | 61.26 | 52.70 | 50.49 | 49.31 | 53.48 | 59.25 | |
@@
Finding 1: PSFT with warm-up consistently outperforms standard SFT on in-domain tasks. Table 1 presents the performance across various in-domain benchmarks, where all fine-tuned methods achieve notable improvements over the Base model. PSFT with the warm-up phase achieves the best overall results, maintaining robustness across different in-domain benchmarks.
@@
Finding 2: PSFT demonstrates strong generalization ability. Table 1 also shows different performance on out-of-domain benchmarks. We find that injecting long CoT data into the model improves its general reasoning ability, consistent with the findings reported in Zhou et al. (2025). In contrast, PSFT exhibits a significantly better generalization ability compared to SFT. For example, in IFEval, which evaluates instruction-following capability, SFT, SFT-KL, and PSFT with warm-up significantly degrade the original results. In contrast, PSFT not only maintains strong in-domain performance but also significantly improves reasoning abilities on diverse out-of-domain tasks, with only minimal compromise in instruction compliance.
@@
Finding 3: PSFT demonstrates robustness across models. For example, when evaluating GPQA and TruthfulQA tasks, SFT and SFT-KL exhibit varying behaviors across different base models—one sometimes compromises performance while the other may boost it. In contrast, PSFT consistently improves performance on both Qwen and Llama models.